The sum of three exam scores is 174. The highest score is three times the lowest score. The middle score is equal to the difference between the highest and lowest scores. Please help me find the middle exam score. I was very confused on trying to set up the problem.
Answer by JoelSchwartz(130)

h=highest score
L=lowest score
m=middle score
h=3L
m=h-L
m=3L-L
m=2L
h+m+L=174
3L+2L+L=174
6L=174
L=29
29*6=174
120+54=174
m=2*29
m=58
h=58+29
h=87
87+58+29=174